x=90-x-36
x+x=90-36
2x=54 2/2x=x 2/54=14
x=14 90-14=76

Answer by johnherbert@yahoo.com(1)

x=90-x-36
x+x=90-36
2x=54 2/2x=x 2/54=14
x=14 90-14=76